1) The following table gives information on the amount of sugar (in grams) and the calorie count in one serving of a sample of 1313 varieties of Kellogg's cereal.

Sugar (grams)

5

15

12

11

8

6

7

3

8

14

20

4

13

Calories

130

200

160

110

120

80

190

120

120

190

190

130

120

Find the predictive regression equation of the number of calories on the amount of sugar.
Carry out all calculations exactly, and round the final answers to three decimal places.

(This is a multiple choice problem, please pick the correct answer out of the following.)

^y=97.932+4.658x
^y=97.932−4.658x
^y=4.658+97.932x
^y=−188.222+5.658x
^y=5.658+188.222x

2) The following table gives information on the amount of sugar (in grams) and the calorie count in one serving of a sample of 1313 varieties of Kellogg's cereal.

Sugar (grams)

6

15

12

11

8

6 7 4 9 14 20 3 13

Calories

120

160

110 140 80 210 100 120 190 190 130 120

The predictive regression equation of the number of calories on the amount of sugar is ^y=95.676+4.892xy^=95.676+4.892x, where xx is amount of sugar (in grams) and yy is calories.

Give a brief interpretation of the values of a=95.676 and b=4.892 in the equation. (this is also M/C, please pick the correct answer for a and b).

The value of a=95.676

means that, on average, the # of calories increases by one for every 95.676 gram increase in sugar?

means that, on average, the # of calories increases by 95.676 for every one gram increase in sugar?

represents the number of calories when the amount of sugar is zero?

The value of b=4.892

means that, on average, the # of calories increases by one for every 4.892 gram increase in sugar?

represents the number of calories when the amount of sugar is zero?

means that, on average, the number of calories increases by 4.892 for every one gram increase in sugar?
Answer #1

1. Y.My (X-M.)2 22.0178 X - My -4.6923 5.3877 2.3077 1.3077 -1.6923 -3.6923 -2.6923 -6.6923 -1.6923 4.3077 10.3077 -5.6923 3.307

Sum of X = 126
Sum of Y = 1860
Mean X = 9.6923
Mean Y = 143.0769
Sum of squares (SSX) = 296.7692
Sum of products (SP) = 1382.3077

Regression Equation = ŷ = bX + a

b = SP/SSX = 1382.31/296.77 = 4.658

a = MY - bMX = 143.08 - (4.66*9.69) = 97.932

ŷ = 4.658X + 97.932

So correct answer here is ^y=97.932+4.658x